How many grams of AlCl3 will be produced if 2.50 moles of Al react?


1
Expert's answer
2021-02-09T04:43:59-0500

At first, the chemical reaction should be written:

2Al + 3Cl2 = 2AlCl3

The coefficients before Al and AlCl3 are equal, so from 2.50 moles of Al 2.50 moles of AlCl3 are obtained.

m=n*M

M(AlCl3)=27+3*35.5=133.5(g/mol)

m(AlCl3)=2.50*133.5=333.75(g).